OpenStudy (anonymous):

Find all the zeros of the equation: -3x^4+27x^2+1200=0 PLEASE HELP!! I don't know how to do this and it would be very helpful if someone could explain

OpenStudy (anonymous):

their are 3 0's

OpenStudy (michele_laino):

first step: I introduce a new variable \(z\) such that \(z=x^2\) so I can reqrite the equation like below: \[ - 3{z^2} + 27z + 1200 = 0\] which can be solved easily with respect to \(z\)

OpenStudy (michele_laino):

oops.. I can rewrite*...

OpenStudy (anonymous):

Okay, so then what is the next step?

OpenStudy (michele_laino):

we have to solve such equation for \(z\). In order to do that, I factor out \(-3\) so I can write this: \[ - 3\left( {{z^2} - 9z - 400} \right) = 0\] tefore we have to solve this equation: \[{z^2} - 9z - 400 = 0\] using the standard formula: \[z = \frac{{ - b \pm \sqrt {{b^2} - 4ac} }}{{2a}}\] where \(a=1,\;b=-9,\;c=-400\)

OpenStudy (michele_laino):

oops.. therefore*

OpenStudy (anonymous):

Okay, give me a min. I'll do that.

OpenStudy (anonymous):

Z=25, or Z=-16?

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

Okay, now what?

OpenStudy (michele_laino):

next I use the definition of the variable \(z\), so I get the two subsequent quadratic equations: \[\begin{gathered} {x^2} = 25 \hfill \\ {x^2} = - 16 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

since the square of any real number is always a positive number, then we can see that the second quadratic equation has not solution

OpenStudy (michele_laino):

namely there is not a number \(n\) such that \(n^2=-16\), am i right?

OpenStudy (anonymous):

Yes.

OpenStudy (michele_laino):

ok! Therefore we have to solve this equation: \(\huge x^2=25\) please what are the solution of such equation?

OpenStudy (anonymous):

x=5

OpenStudy (michele_laino):

hint: what is \((-5)^2=...?\)

OpenStudy (anonymous):

25

OpenStudy (michele_laino):

thta's right! So the real solution of your starting equation, are: \(\huge x_1=5,\;x_2=-5\)

OpenStudy (michele_laino):

oops..solutions*

OpenStudy (anonymous):

Okay! Now what?

OpenStudy (michele_laino):

we have finished. More precisely, your equation is a fourth grade equation, so we have two real solutions, and two imaginary solutions

OpenStudy (anonymous):

So it's 5, -5, and then what?

OpenStudy (michele_laino):

the other imaginary solutions, are given solving this equation: \(x^2=-16\) such solutions are: \(x_3=4i,\;x_4=-4i\) where \(i\) is such that \(i^2=-1\)

OpenStudy (anonymous):

Okay! That actually makes sense. Thanks so much for the help! You're a lifesaver.

OpenStudy (michele_laino):

:)